OpenStudy (anonymous):

Find the sum or difference. 16/17+3/17 A 19/17 B 9/34 C 1/2

OpenStudy (freckles):

you have the same denominator just add the tops

OpenStudy (freckles):

16+3=?

OpenStudy (freckles):

\[\frac{16}{17}+\frac{3}{17}=\frac{16+3}{17}=\frac{?}{17}\]

OpenStudy (anonymous):

so 19/17

OpenStudy (freckles):

yep